OpenStudy (anonymous):

What will it cost to carpet a room with indoor/outdoor carpet if the room is 10 feet wide and 12 feet long? The carpet costs 12.51 per square yard. A. $166.80 B. $175.90 C. $184.30 D. $189.90 E. $192.20

Directrix (directrix):

I assume the room being carpeted is rectangular shaped. The "catch" here is that the carpet dimensions are in feet (area in square feet) and the cost in square yards.

OpenStudy (anonymous):

easiest to convert feet to yard, find the area and multiply by cost/yard^2

Directrix (directrix):

Ten feet is 10/3 yards and twelve feet is 12/3 yards. The area of the carpet is the area of a rectangle. Area = (10/3) * (12/3) = 120/9 square yards which simplifies to 40/3 square yards. The carpet costs 12.51 per square yard. $12.51 * ( 40/3) = $166.80

Directrix (directrix):

@Pluto113 --> Question?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Does anyone know about above?

OpenStudy (anonymous):

A